A line passing through (6, 9) and (5, y) is parallel to a line with slope 5. What is the value of y?
Found 2 solutions by scott8148, edjones:
Answer by scott8148(6628)

parallel means the same slope

slope of 5 means y changes 5 times as fast as x

x decreases by 1 so y decreases by 5 ... y=9-5 ... y=4

Answer by edjones(8007)

slope=m=change in y/change in x
5=(9-y)/(6-5)
5=9-y (6-5)=1
-y=-4 subtract 9 from each side.
y=4 multiply -1 times each side.
Ed